Solve for x;
5(8-x) = 3x

Answer by bucky(2189)

5(8-x) = 3x
.
do the multiplication on the left side by multiplying 5 times each of the terms in the parentheses
to get:
.
5*8 - 5*x = 3x
.
This simplifies to:
.
40 - 5x = 3x
.
add 5x to both sides to eliminate the -5x from the left side. This gets the term with x
on one side of the equal sign and the number term on the other side:
.
40 = 8x
.
Now divide both sides by the multiplier of x (which is 8) to find x. When you divide
both sides by 8 the equation becomes:
.
40/8 = 8x/8
.
which reduces to:
.
5 = x
.
The solution to this problem is x = 5
.
Hope this helps you to see a logical pattern of thought in solving linear algebraic
equations.